# Offline Imitation of Badminton Player Behavior via Experiential Contexts and Brownian Motion

### v1

@@ IV-B Quantitative Results (RQ1)
Table I presents the performance of RallyNet and the baselines. We summarize the observations as follows:
IV-B1 Superior Performance of RallyNet
@@
Clearly, RallyNet surpasses all the baselines in terms of all metrics, whether given only the initial state, or the state of the first two steps. Quantitatively, RallyNet outperforms all the baselines by at least 7.08 in terms of CTC loss when predicting the shot types, and achieves lower DTW distances than the baselines by at least 8.27% and 1.16% in the predicted landing and moving positions, respectively. Note that since players usually have similar positions and shot types for serving and receiving (i.e., the first two steps), the prediction after two steps is more difficult to predict. These results demonstrate that RallyNet addresses the compounding error in offline imitation learning for turn-based sports well.
IV-B2 Hierarchical BC Advantage
@@
We observe that IQ-Learn produces unfavorable performance, possibly due to the difficulty of inferring the reward function in a scenario where the opponent exclusively determines the next state. This may lead to a weak correlation between the current and the next state, making IQ-Learn unsuitable for this particular task. HBC can better capture complex player strategies as it can switch between different strategies by selecting different options based on the different states. We can observe that HBC achieves better prediction than BC, and manifests the importance of using HIL in the turn-based setting.
IV-B3 The Importance of Leveraging Experience
@@
Existing supervised learning methods for turn-based sports, such as ShuttleNet and DyMF, diverge from imitation learning models. These models, structured on sequence-to-sequence architectures, predict player actions by relying on historical rally data. DyMF excels in predicting shot types, while ShuttleNet performs well in forecasting player movements, showcasing its proficiency in the handling of player interactions and rally dynamics. However, these approaches face challenges as errors accumulate during extended rallies. In contrast, RallyNet stands out by leveraging experience, enabling it to emulate player-like behavior even in prolonged rallies. This distinctive capability serves to alleviate the cumulative errors observed in other methods, resulting in more realistic and accurate behavioral predictions.
IV-C Length Distribution Difference (RQ2)
@@
The length of generated rallies serves as a reflective measure of whether the model has learned realistic gameplay situations. A model capable of capturing rally duration can reflect a genuine simulation of the match. To visually depict the distribution of generated rally lengths in comparison to real rally lengths, we utilize Kernel Density Estimate (KDE). Additionally, the quantitative dissimilarity between these two distributions is evaluated using the Jensen–Shannon divergence (JSD). In Figure 6, BC struggles to effectively capture player interactions, resulting in premature or delayed termination. In contrast, both HBC and RallyNet, by modeling more intricate strategies, produce generated rally length distributions close to the real ones. Notably, even a rule-based agent using our carefully designed experience extraction function can achieve a similar rally length distribution, emphasizing that our experience extraction captures rally duration to aid RallyNet in making more realistic decisions. Figure 7 depicts the generated length distributions of ShuttleNet and DyMF alongside real ones. It is evident that ShuttleNet and DyMF, not being imitation learning models but sequence-to-sequence tasks, face the challenge in generating ideal rally lengths. This challenge stems from the infrequent occurrence of error behaviors in dataset, where a rally involves only one error behavior (e.g., shuttlecock out of bounds or hitting the net), posing a significant obstacle for these models to accurately capture rally duration.
@@
Figure 7: The length distribution difference between generated rallies and ground truth rallies. It is evident that existing sequence-to-sequence methods struggle to capture the timing of player error behavior.
IV-D Win Rate Difference (RQ3)
To evaluate the outcome consistency of learned models in mimicking real player behavior, we adopt a straightforward approach by examining win rates among players. As depicted in Table II, considering the diverse quantities of match data available for each player in the dataset, our analysis focuses on two real-world professional players, X and Y, who possess the most extensive match records. For both X and Y, we chose four opponents from their respective match histories, subsequently comparing the generated rally win rates with the simulated win rates from the models.
To ensure the validity of win rate observations within realistic gameplay scenarios, we evaluated the performance of the top three imitation learning models: BC, HBC, and RallyNet. Table II illustrates that RallyNet consistently demonstrates smaller differences from the actual win rate across a majority of matchups, with a maximum difference of merely 8.15% and an average difference of 3.79%. In contrast, BC and HBC exhibit differences of up to 22.5% in specific matchup combinations, with an average difference surpassing RallyNet by at least 7.16%. These findings not only underscore RallyNet’s stability but also highlight its exceptional capability to accurately replicate real match outcomes.
TABLE II: Win rate differences between ground truths and learned agents. X and Y denote two actual players chosen from the dataset. A, B, C, D are the four opponents of player X, while E, F, G, H are the four opponents of player Y, selected from their past matches. Lower differences indicate better alignment between the observed win rates and those predicted by the learned agents.
@@ IV-D Win Rate Difference (RQ3)
TABLE III: Ablation study of our model.
Given initial state only Given states of first two steps
BC 0.9603 63.5967 0.4829 1.1199 1.0343 56.1007 0.5832 1.3084
BC w/ ECS 0.6877 27.0127 0.4577 1.6976 0.9259 29.4051 0.5780 2.0416
Indiv. RallyNet 0.6063 19.0623 0.3766 1.9404 0.8189 21.3749 0.5305 2.4782
IV-E Ablation Studies (RQ4)
As RallyNet can be seen as an extension of BC with the addition of ECS and LGBM, an ablation study was conducted by developing two variants to investigate the relative contributions of two components: 1) BC w/ ECS, which is RallyNet without the LGBM module, 2) Indiv. RallyNet, which is RallyNet with the proposed LGBM replaced by the individual LGBMs (i.e., two independent stochastic processes, one for each player). The result is shown in Table III.
IV-E1 The Effect of ECS
Recall that ECS was designed to capture the agent’s intent through experience, enabling the agent’s behavior throughout the rally to not be influenced by partially incorrect decisions. The agent can already achieve a moderately low prediction error by adding ECS to BC, especially when predicting the shot type and the landing position. However, without the help of LGBM, the agent suffers from a large error in moving position since it disregards the opponent’s behavior.
IV-E2 The Effect of LGBM
Recall that LGBM is meant to jointly capture the interaction of players to help the agent generate more realistic behavior. RallyNet (i.e., BC w/ ECS w/ LGBM) demonstrates a significant improvement in performance, particularly in the prediction of moving positions, compared to the model BC w/ ECS. Furthermore, we investigated the effect of replacing the proposed LGBM with individual LGBMs (i.e., Indiv. RallyNet). The results showed that without the shared inductive bias introduced by proposed LGBM, the agent ignored players’ interaction, resulting in a degradation of the predicted moving position performance. Overall, the results further support the effectiveness of LGBM in RallyNet and provide evidence that RallyNet captures the alternative decision-making nature of the turn-based sports.
IV-F Sensitivity Analysis (RQ5)
IV-F1 Loss weights
The loss that we are minimizing is:
$\mathcal{L}=w_{1}\cdot\mathcal{L}_{pred}+w_{2}\cdot\mathcal{L}_{ctx}+w_{3}\cdot\mathcal{L}_{sde},$ (20)
where $w_{1}$, $w_{2}$, $w_{3}$in [0, 1] are hyper-parameters to balance the weights of the corresponding losses. We have set the loss weights $w_{1},w_{2},w_{3}$ to 1. Considering the significant scale difference between $\mathcal{L}_{land}$, $\mathcal{L}_{move}$, and $L_{reg}$ compared to $\mathcal{L}_{type}$, we scale down $\mathcal{L}_{land}$, $\mathcal{L}_{move}$, and $\mathcal{L}_{reg}$ by a factor of 0.01. The purpose of $\mathcal{L}_{reg}$ is to add a regularization loss that prevents overlap between the predicted landing distributions and moving distributions, ensuring that the model does not degenerate into a simple policy. This regularization loss is defined as the average negative distance between bivariate normal distributions. The relationship between $\mathcal{L}_{type}$, $\mathcal{L}_{land}$, $\mathcal{L}_{move}$, and $\mathcal{L}_{reg}$ is given by:
$\mathcal{L}_{pred}=\mathcal{L}_{type}+\frac{(1-\alpha)}{100}\cdot\mathcal{L}_{(land,move)}+\frac{\alpha}{100}\cdot\mathcal{L}_{reg},$ (21)
where the regularization loss weight $\alpha$ is set to 0.05 for all experiments. To investigate the influence of varying loss weights on the model’s performance, we conducted a parameter analysis by examining different values for $w_{1}$, $w_{2}$, and $w_{3}$. Specifically, we evaluated the model’s MRNS using weights of 0.25, 0.50, and 0.75 for each loss term. The results of these experiments are presented in Figure 8.
The results indicate that increasing the weights of $\mathcal{L}_{pred}$ and $\mathcal{L}_{ctx}$ improves performance by enhancing the accuracy of target process outputs and context representation learning. This enables the player process used for inference to closely approximate the precise target process. We also observed that the influence of $\mathcal{L}_{sde}$ on performance is stable. This suggests that when the target process is effectively learned, the player process effortlessly approximates the target process, resulting in consistent performance that is minimally affected by $\mathcal{L}_{sde}$.
Figure 8: MRNS Performance of RallyNet w.r.t different value of loss weights.
IV-F2 Context dimension
In all experiments, we set the context embedding dimension to 128. The dimension of the context embedding plays a crucial role in determining the continuity of the context space and influencing the specificity of the selected context (i.e., player’s intention). As shown in Figure 10, increasing the dimension of the context embedding results in improved overall performance.
IV-F3 The number of bivariate Gaussian distributions
Across all experiments, we utilized a fixed number of 5 bivariate Gaussian distributions for predicting landing positions and moving positions. The number of distributions played a crucial role in the accuracy of these predictions. We conducted experiments with three different numbers of bivariate Gaussian distributions: 2, 5, and 8. Figure 10 reveals that insufficient or excessive distributions were unable to adequately capture the intricate distributions of landing positions and moving positions.
Figure 9: MRNS Performance of RallyNet w.r.t different value of context dimensions.
Figure 9: MRNS Performance of RallyNet w.r.t different value of context dimensions.
Figure 10: MRNS Performance of RallyNet w.r.t different number of distributions
IV-F4 Cross-Domain Evaluation
We cannot test our model on other turn-based sports since there is only one existing public dataset that can be accessed. To further illustrate the effectiveness of RallyNet, we have included additional experiments by comparing RallyNet with our baselines in the Pong environment of Atari games, which shares some turn-based characteristics. The state of the Pong environment is an image, and the action is a discrete one that includes left and right movement. To train an expert, we used DQN and captured the position of each shot of the agent and the computer opponent simultaneously. We then converted the data into those of the same turn-based data format as our badminton dataset, including landing position, moving position, and whether the ball is reachable. We also rescaled the court to between -1 and 1. Regarding RallyNet on Pong, we used the same implementation as that for badminton. It is worth mentioning that ShuttleNet and DyMF have the limitation wherein they require information of the previous two steps to make predictions. Thus, they cannot be directly compared to RallyNet and the baselines of imitating rally-wise player behavior. Table IV illustrates the comparative results of imitating player behavior in the Pong environment. The results show that RallyNet outperforms all the baselines and indicate that RallyNet can address the compounding error in offline imitation learning for other turn-based sports environments.
TABLE IV: Quantitative results on the Pong environment of Atari games.
Given initial state only Given states of first two steps
Model Land Shot Move Land Shot Move
BC 0.4025 5.2086 0.5764 0.4859 1.5431 0.5618
HBC 0.4079 4.5921 0.4787 0.4960 0.6891 0.5162
RallyNet 0.3884 4.4488 0.4495 0.4513 0.5738 0.4764
IV-G Case Studies (RQ6)
IV-G1 Simulation of Player Behavior
We describe a use case of the learned model in sports analytics for characterizing the style of a professional player by simulating the behavior of the player (e.g., the landing and the subsequent moving position upon a shot) against different opponent players in different scenarios. Figure 11 shows the simulated distributions of the landing and moving position of a player after a defensive shot. In this example: (i) This player tends to have a landing position mostly on both sides of the backcourt and only occasionally on the midcourt. (ii) This player tends to move to the midcourt after a defensive shot to prepare for the next stroke.
@@ IV-G1 Simulation of Player Behavior
To provide a more comprehensive view of the results, we have generated visualizations comparing the landing distributions of RallyNet and BC across various states. Figure 12 illustrates the simulated landing distributions of under different state conditions. It demonstrates that RallyNet consistently captures players’ decision patterns with precision. In contrast, BC exhibits instances where it manages to approximate the distribution roughly but with deviations, and occasionally even exhibits significant misjudgments. Accordingly, such characterization can help the coach better understand the style of the player and thereby devise tactical plans.
Figure 11: The landing distributions and the moving distributions of a player when using a defensive shot.
Figure 12: The landing distributions of the player in various states are shown, with RallyNet in the upper row and BC in the lower row. RallyNet excels in capturing landing distribution compared to BC. Note that these visualizations portray results extracted from the testing dataset corresponding to each specific state, not the model’s predicted landing distribution for a specific state.
IV-G2 Tactical Interpretation of Player Behavior
RallyNet strategically interprets player behavior by selecting a context as the agent’s intent, guiding its actions accordingly. The intent is decoded by the context decoder of the ECS, revealing the expected action sequence. Figure 13 exemplifies this process, deciphering the intent behind a shot based on the preceding (long serve in the green grid) and current shots (net shot in the yellow grid). For instance, if Player A’s intent is decoded as planning a subsequent smash by strategically positioning the opponent in the backcourt, it suggests a tactical move to set up aggressive shot returns. This showcases RallyNet’s capability to provide insightful tactical interpretations, a valuable asset for badminton coaches and players.
To validate the system’s ability to decode the physical meaning of selected contexts, we conducted a reconstruction experiment. This evaluated the reconstruction performance of the context decoder in ECS, showcasing its effectiveness in terms of capturing the player’s intent from their actions. Unlike focusing on player interactions, this approach directly extracts the underlying intent from the action sequence. Table V demonstrates that the reconstruction performance surpasses the imitation performance, showcasing the effectiveness of our context encoder in capturing and decoding player intent.
@@ IV-G2 Tactical Interpretation of Player Behavior
Figure 13: An example of interpreting the intent behind a player’s action. The clear stroke by player A is meant to plan for a subsequent smash.
TABLE V: The reconstruction performance of the context encoder.
Model Land($\downarrow$) Shot($\downarrow$) Move($\downarrow$)
RallyNet (Ours) 0.4011 4.1918 0.2009
V Related Work
V-A Inverse Reinforcement Learning
Inverse Reinforcement Learning (IRL) [36, 37, 38] is an IL approach that attempts to learn the underlying rewards that an expert is optimizing from the demonstrations. With the learned reward function, IRL can guide the agent’s behavior toward that of the expert. However, in turn-based sports, each player’s actions determine the next state for other players. Treating opponents as part of the environment is equivalent to training in a constantly changing environment, which can make it challenging for IRL to learn optimal policies.
V-B Offline Imitation Learning via Behavior Cloning
Behavior cloning (BC) provides a form of supervised learning for training policies by learning direct mapping from states to actions, which can be used in the offline setting [28, 39]. Recent works such as DT [40], GCSL [41], and RvS [42] have shown that not only is supervised learning and bypassing the learning reward function able to attain better results for offline learning, but it is also easier to use and is more stable than offline inverse reinforcement learning (IRL) [36, 43, 27]. Moreover, previous work [44, 29] has extended the BC to hierarchical BC by dividing a task into several sub-tasks, where the low-level behaviors are controlled by high-level decisions to capture long-term decision-making processes. However, these existing approaches focus on the complex behavior and the long-term task of the same target, but neglect the characteristics of the mixed sequences, which causes serious compounding errors.
V-C Badminton Tactical Analysis
AI has significantly contributed to the domain of badminton sports analytics, addressing diverse challenges. For instance, [45] introduced the Badminton Language for Sequence Representation (BLSR), presenting a unified language capable of efficiently converting match videos into analyzable datasets. BLSR, characterized by its human-readable, universal, and professional nature, enables the straightforward interpretation of entire match processes without the need for video footage. In strokes forecasting, [9] advanced the prediction of players’ future strokes based on their past actions. The innovative ShuttleNet framework incorporates a position-aware fusion of rally progress and player styles, accounting for dependencies at each stroke. Additionally, [25] addressed the movement forecasting task, encompassing stroke and player movement predictions. They employed DyMF, a novel model based on dynamic graphs and hierarchical fusion grounded in player movement graphs. However, past models focused on predicting future actions given specific initial steps, but the lack of realistic agents hampers accurate match simulation. We propose a closed-loop AI approach [1]. This system includes six key components to enhance badminton players’ performance: Badminton Data Acquisition, Imitating Players’ Style, Simulating Matches, Optimizing Strategies, Training Execution, and Real-World Competitions. In this paper, we only focus on one specific task within the closed-loop AI approach: imitating players’ styles. We propose RallyNet to imitate badminton players’ styles. Beyond evaluating RallyNet using previously proposed offline metrics [1], we assess the learned agent’s applicability to different badminton scenarios, confirming the feasibility of RallyNet in badminton games simulation.
VI Conclusion
@@
In this work, we present RallyNet, a novel hierarchical offline imitation learning model for learning player decision-making strategies in turn-based sports. By modeling players’ decision-making processes as CMDP, we introduce ECS which leverages experiences to construct a context space, and selects the context as the agent’s intent. This reduces the impact of partial decision errors on overall behavior. The ECS component reduces decision-making errors of agents in the CMDP setting, even if it is not turn-based, while LGBM captures player interactions to generate more realistic behaviors for various turn-based sports. We believe that the innovative ideas, connecting player intents with interaction models using GBM and offering an understanding of real interactions in turn-based sports, have substantial potential to inspire researchers across other turn-based sports. Our evaluation of a real-world badminton dataset shows that RallyNet outperforms existing offline IL and the state-of-the-art turn-based supervised method, and has potential for sports analytics. Furthermore, we demonstrate that RallyNet is indeed a promising solution to sports analytics via practical use cases.

### v2

@@ 5.2 Quantitative Results (RQ1)
Table 1 presents the performance of RallyNet and the baselines. We summarize the observations as follows:
@@
Superior Performance of RallyNet. RallyNet surpasses all the baselines in terms of all metrics, whether given only the initial state, or the state of the first two steps. Quantitatively, RallyNet outperforms all the baselines by at least 7.08 in terms of CTC loss when predicting the shot types, and achieves lower DTW distances than the baselines by at least 8.27% and 1.16% in the predicted landing and moving positions, respectively. Note that since players usually have similar positions and shot types for serving and receiving (i.e., the first two steps), the prediction after two steps is more difficult. These findings highlight RallyNet’s effective mitigation of compounding errors in badminton player imitation.
@@
Hierarchical BC Advantage. We observe that IQ-Learn produces unfavorable performance, due to the difficulty of inferring the reward function in a scenario where the opponent exclusively determines the next state. This may lead to a weak correlation between the current and the next state, making this IRL method unsuitable for this particular task. HBC can better capture complex behavior as it can switch between different strategies by selecting different options based on varying states, highlighting the importance of using HIL in turn-based sports.
@@
The Importance of Leveraging Experience. Existing supervised learning methods for turn-based sports, such as ShuttleNet and DyMF, diverge from imitation learning models. These models, structured on sequence-to-sequence architectures, predict player actions by relying on historical rally data. DyMF excels in predicting shot types, while ShuttleNet performs well in forecasting player movements, showcasing its proficiency in the handling of player interactions and rally dynamics. However, these approaches face challenges as errors accumulate during prolonged rallies. In contrast, RallyNet stands out by leveraging experience, enabling it to emulate player-like behavior even in long rallies.
5.3 Length Distribution Difference (RQ2)
@@
The length of generated rallies serves as a reflective measure of whether the model has learned realistic gameplay situations. To visually depict the distribution of generated rally lengths in comparison to real rally lengths, we utilize Kernel Density Estimate (KDE). Additionally, the quantitative dissimilarity between these two distributions is evaluated using the Jensen–Shannon divergence (JSD). In Fig. 3, BC struggles to effectively capture player interactions, resulting in premature or delayed termination. In contrast, both HBC and RallyNet, by modeling more intricate strategies, produce generated rally length distributions close to the real ones. Notably, even a rule-based agent using our carefully designed experience extraction function can achieve a similar rally length distribution, emphasizing that our experience extraction captures rally duration to aid RallyNet in making more realistic decisions. However, models like ShuttleNet and DyMF, being sequence-to-sequence models rather than imitation learning models, face challenges in generating ideal rally lengths. This challenge stems from the infrequent occurrence of error behaviors in dataset, where a rally involves only one error behavior (e.g., shuttlecock out of bounds or hitting the net) , making it difficult for these models to accurately capture rally duration.
@@
5.4 Win Rate Difference (RQ3)
To evaluate the outcome consistency of learned models in mimicking real player behavior, we adopt a straightforward approach by examining win rates among players. As depicted in Table 2, considering the diverse quantities of match data available for each player in the dataset, our analysis focuses on two real-world professional players, X and Y, who possess the most extensive match records. For both X and Y, we chose four opponents from their respective match histories, subsequently comparing the generated rally win rates with the ground truth win rates. To ensure the validity of win rate observations within realistic gameplay scenarios, we evaluated the performance of the top three imitation learning models: BC, HBC, and RallyNet. Table 2 illustrates that RallyNet consistently demonstrates smaller differences from the actual win rate across a majority of matchups, with a maximum difference of merely 8.15% and an average difference of 3.79%. In contrast, BC and HBC exhibit differences of up to 22.5% in specific matchup combinations, with an average difference surpassing RallyNet by at least 7.16%. These findings not only underscore RallyNet’s stability but also highlight its exceptional capability to accurately replicate real match outcomes.
Table 2: Win rate differences between ground truths and learned agents. X and Y denote two actual players chosen from the dataset. A, B, C, D are the four opponents of player X, while E, F, G, H are the four opponents of player Y. Lower differences indicate better alignment between observed and predicted win rates by the agents.
@@ 5.4 Win Rate Difference (RQ3)
5.5 Case Studies (RQ4)
@@
Simulation of Player Behavior. We describe a use case of the learned model in sports analytics for characterizing the style of a professional player by simulating the behavior of the player in different conditions. We generated visualizations comparing the landing distributions of RallyNet and BC across various states. Fig. 5 illustrates the simulated landing distributions of under different state conditions. It demonstrates that RallyNet consistently captures players’ decision patterns with precision. In contrast, BC exhibits instances where it approximates the distribution roughly but with large deviations, and occasionally even exhibits significant misjudgments. Accordingly, such characterization can help the coach understand the style of the player and thereby devise tactics.
Figure 4: The landing distributions of the player in various states are shown, with RallyNet in the upper row and BC in the lower row.
@@
Figure 5: Interpreting the intent behind an action. player A’s clear stroke is meant to plan for a smash.
Tactical Interpretation of Player Behavior. RallyNet strategically interprets player behavior by selecting a context as the agent’s intent, guiding its actions accordingly. The intent can be decoded by the context decoder of the ECS, revealing the agent’s expected action sequence. Fig. 5 illustrates an example to investigate the underlying intent behind a shot based on the previous shot (long serve in the green grid) and the current shot (net shot in the yellow grid). For instance, if Player A’s intent is decoded as planning a subsequent smash by strategically positioning the opponent in the backcourt, it suggests a tactical move to set up aggressive shot returns. This showcases RallyNet’s capability to provide insightful tactical interpretations, a valuable asset for badminton coaches and players.
6 Conclusion
@@
In this work, we present RallyNet, a novel hierarchical offline imitation learning model for learning player decision-making strategies in turn-based sports. By modeling players’ decision-making processes as CMDP, our ECS leverages experiences to construct a context space and selects the context as the agent’s intent, allowing the reduction of decision-making errors in agents’ overall behaviors. Meanwhile, LGBM captures player interactions to generate more realistic behaviors. Extensive evaluations on real-world badminton singles matches show that RallyNet outperforms existing offline IL as well as the state-of-the-art turn-based supervised method with multifactorial assessments, and showcases the practicability for sports analytics. We believe RallyNet serves as a general framework for other turn-based sports due to the flexible design for modeling player intents and offering an understanding of real interactions.